The shape of the distribution of the time required to get an oil change at a 20-minute facility is unknown. however, records indicate records indicate that the mean time is 21.4 minutes and the standard deviation 4.9 minutes.

B) what is the probability that a random sample of n=40 oil changes results in a sample mean time less than 20 minutes?

the probability is approximately _____
Answer #1

Solution :

Given that ,

mean = \mu = 21.4

standard deviation = \sigma = 4.9

n = 40

\mu\bar x = 21.4

\sigma\bar x =\sigma  / \sqrt n = 4.9/ \sqrt 40=0.7748

P(\bar x <20 ) = P[(\bar x - \mu \bar x ) / \sigma \bar x < (20 - 21.4) /0.7748 ]

= P(z <-1.81 )

Using z table

= 0.0351